OpenStudy (anonymous):

The dimensions of a rectangle are such that its length is 5 inches more than its width. If the length were doubled and if the width were decreased by 2 inches, the area would be increased by 136 in^2. what are the length and width of the rectangle?

OpenStudy (anonymous):

l = w+5 (2l)(w-2) = lw + 136 Substitute for l. Solve quadratic in w. Then go from there.

OpenStudy (anonymous):

L = 5 + W 2L(W-2) = LW + 136 2(W + 5)(W-2) = (W + 5)W + 136 2w^2 + 6w - 20 = w^2 + 5w + 136 w^2 + w -156 = 0 Factors into (w - 12)(w + 13) Solutions are 12 and -13. You can't have a negative length so that means the width is 12. L = 5 + w L = 5 + 12 L = 17, W = 12